# Evaluating Meta-Regression Techniques: A Simulation Study on Heterogeneity in Location and Time

### v1

@@ Abstract
In this paper, we conduct a simulation study to evaluate conventional meta-regression approaches (study-level random, fixed, and mixed effects) against seven methodology specifications new to meta-regressions that control joint heterogeneity in location and time (including a new one that we introduce). We systematically vary heterogeneity levels to assess statistical power, estimator bias and model robustness for each methodology specification. This assessment focuses on three aspects: performance under joint heterogeneity in location and time, the effectiveness of our proposed settings incorporating location fixed effects and study-level fixed effects with a time trend, as well as guidelines for model selection. The results show that jointly modeling heterogeneity when heterogeneity is in both dimensions improves performance compared to modeling only one type of heterogeneity.
@@ 1 Introduction
A meta-analysis is a useful tool that combines results from independent studies with similar design to examine whether there is consistency in the results across studies or if those results conflict (McGrath et al., 2019; Bakbergenuly and Kulinskaya, 2018; Kelley, 2012; Rice et al., 2018). Note that meta-regressions are commonly referred to as meta-analyses (and vice versa), since the only difference introduced by a meta-regression is that it transcends the univariate model of a meta-analysis by including identifiable study and/or subject characteristics as covariates. Data from the different studies under consideration can be at the subject level or aggregate level using the summary statistics of each study; this manuscript will focus on the subject level due to the many issues inherent in aggregate level data (Cooper & Patall, 2009; Stewart & Clarke, 1995; Lambert et al., 2002; Yamaguchi et al., 2014). Consequently, thousands of meta-analyses have been performed, primarily in the medical, pharmaceutical, ecology, and social sciences literature (Borenstein et al., 2009). The methodology used for meta-analyses is designed to accurately estimate the effect sizes and mitigate any heterogeneity that may be present between each of the studies in the meta-analysis. The standard methodology uses study-level effects: most commonly random effects, fixed effects, and on rare occasions mixed effects (Gambarota & Altoè, 2024; Sugasawa & Noma, 2021; Richardson et al., 2021; Pateras et al., 2018; T. D. Stanley & Doucouliagos, 2017; T. Stanley et al., 2023). But what about heterogeneity in location or time? It is common for meta-analyses to include decades worth of studies from many locations around the world, which introduces the likely variation in both the time and location in which each study was conducted (Thöni & Volk, 2021; Van Den Akker et al., 2020; Engel, 2011; Oosterbeek et al., 2004). In the experimental economic and meta-analysis simulation literature, almost all meta-analyses are conducted at the study level (dominantly using random effects), with only a small percentage controlling for location variation and none controlling for time variation. [^1] Table 4 in Appendix A shows a representative sample of meta-regressions in experimental economics and how they control for heterogeneity: only about 30% control for location heterogeneity, only about 15% control for precise country-level heterogeneity, and only about 5% control for time heterogeneity (implicitly). This manuscript tests three hypotheses when the data follows the standard assumptions:
@@
This research aims to test which meta-regression methodology performs optimally when joint heterogeneity in the location and time between studies in a meta-regression is present in varying degrees. To achieve our research objective, we conduct a simulation study that tests the performance of ten meta-regression methodology specifications that include the standard methodology and the specifications that control for heterogeneity in location and/or time. Note that most of the methodology specifications are fixed-effect specifications since the assumption of random effects is a phenomenon that rarely holds in practice (which is required for random or mixed effect models, please see Section 5.1 for further details) (Wooldridge, 2010; Mundlak, 1978; Hausman & Taylor, 1981; Baltagi, 2005; Clark & Linzer, 2015). In summary, our simulation found that the standard methodology performed the best and the specifications that control for joint heterogeneity performed a close second best. This excludes study-level fixed effects with a trend proposed in the meta-analysis simulation literature because this provided inaccurate trend estimates, but includes a new specification inspired by the literature and introduced by this manuscript: location-level fixed effects with a trend. Accordingly, specifications that did not jointly control for location and time heterogeneity performed the worst (including location-level fixed effects proposed in the experimental economics literature). Note that these results are for joint heterogeneity, in future research we plan to test when only location-level heterogeneity is present. The organization of this paper is as follows: Section 2 reviews the relevant literature, Section 3 shares the methodologies of a standard meta-analysis, and Section 4 shows the additional methodology specifications that control for location and/or time heterogeneity. The simulation in Section 5 tests the performance of each of the methods in the following way: simulation design, the performance criteria of each specification, and which specification performs optimally. Section 6 shows the utility of example applications as well as how simulations can be extended beyond purely synthetic data by extracting parameters from real data. Lastly, Section 7 provides concluding remarks, modeling recommendations for practitioners, and notes on the related areas we plan to study in the future.
@@ 2 Literature Review
There is dense literature that argues whether meta-analyses and meta-regressions should utilize a study-level fixed effects model or a study-level random effects model (Hunter & Schmidt, 2000; Brockwell & Gordon, 2001; Viechtbauer, 2007; Gambarota & Altoè, 2024; Sugasawa & Noma, 2021). Random effects are the most dominant because their design handles data randomly drawn from a larger population (such as experimental subject data), can capture the sampling error of the participants, and can also capture the study-level variance of the participant sampling for each experiment (Kelley, 2012; Higgins et al., 2009; Cooper & Patall, 2009). However, a random effects model requires that variation in draws of the random effect are uncorrelated with variation in all other explanatory variables, which is rare in practice (Wooldridge, 2010). Another argument for using random effects instead of fixed effects at the study level is that it can model heterogeneity instead of indiscriminately controlling for it by averaging (Gambarota & Altoè, 2024; Richardson et al., 2021; Pateras et al., 2018; Erez et al., 1996). There is also a less common mixed effects model (also known as the generalized linear mixed model). This model combines fixed and random effects to capture both the variation within each group (with fixed effects) as well as the variation across groups (with random effects); mixed effects are commonly used for hierarchical data where data can be categorized into groups across different levels (Papadimitropoulou et al., 2019; Bakbergenuly and Kulinskaya, 2018; Stijnen et al., 2010; Overton, 1998; Viechtbauer, 2010). There are important trade-offs to consider for all of these methodology specifications. For instance, one study found an overestimation of standard errors with random effects and an underestimation of standard errors with fixed effects (Overton, 1998). The tradeoffs are compounded for mixed effects since it includes the benefits of both fixed and random effects, but also the limitations of both fixed and random effects. This is related to our research because this is the standard methodology that is tested and compared to our methodology specifications that control location and/or time. To recap, do not indiscriminately use random effects; it has assumptions and limitations of their own, just like fixed effects and mixed effects.
@@
This literature also covers methodology performance when varying degrees of study-level heterogeneity are present (McGrath et al., 2019; Bakbergenuly and Kulinskaya, 2018; Sánchez-Meca & Marín-Martínez, 2008; Hedges & Vevea, 1998; Biggerstaff and Tweedie, 1997). The common finding is that when higher heterogeneity between studies is present (compared to lower heterogeneity), neither study-level random or fixed effects perform as well. Our research extends this to varying degrees of heterogeneity at the location and time level. This is important since the literature overlooks the possibility of variation in location or time in a meta-analysis (excluding only a select few studies). The only exception are spatiotemporal meta-analyses, but the focus of this field of literature is very different since these use spatial economic models to model the heterogeneity between locations by using the distance between them along with other related covariates (B. T. Johnson et al., 2017). In contrast, when this manuscript describes location heterogeneity, its focus is closer to two way fixed effects in panel data models: distance apart is not a factor, we solely control whether locations are different or not. Note that many of the sources that highlight spatiotemporal heterogeneity is present, don’t account for it explicitly in their modeling (Liu & Xiao, 2022; Ren et al., 2022; Jiang et al., 2024). Recently, some of those studies have explored controlling for location-level or time-level effects more directly than implicitly through study-level effects. In the experimental economic literature, location-level effects were controlled by country, geographic region, or continent (Aimone et al., 2024; Oosterbeek et al., 2004; N. D. Johnson & Mislin, 2011). Our research analyzes the performance of each specification and how well their specification captures varying degrees of location-level heterogeneity, neither of these have been done since the aforementioned meta-analyses are only empiric.
@@
The meta-analysis simulation literature explores time-level effects, but each study controls or measures time differently, so there is no set consensus. Accordingly, our research aims to set a consensus by testing the relevant specifications that control for time heterogeneity. Two studies measure fixed effects with a trend for experiments that take place at different time periods, and another study introduces a meta-analysis methodology called Continuous Time Meta-Analysis (CoTiMA) that includes time lags for both within each study and between the studies inside of a meta-analysis (Dormann et al., 2020; LeBlanc & Banks, 2022; Rudolph et al., 2020). Closest to our research is a study that introduces a meta-analysis methodology to model variation over time between the static studies inside the meta-analysis (LeBlanc & Banks, 2022). This methodology models variation in time by testing if there are any trends in each static study across time by category. So, only the date in the meta-analysis is dynamic; there are no time lags within each study. But, either time-level fixed effects or fixed effects with a trend capture the variation of time between studies (please see Section 4 for further details). Both studies used study-level fixed effects with a trend, so we are going to test this specification in the face of joint location and time heterogeneity and compare its performance to other specifications. These other specifications include the standard methodology, location-level fixed effects with a trend, and others that control for location and/or time. Our manuscript is the first to test the following meta-regression specifications together in a simulation: location-level fixed/random effects, time-level fixed effects, and location-level fixed effects with trend. The goal of testing these methods are different from spatiotemporal heterogeneity meta-analyses because they control for the heterogeneity instead of spatially modeling it. The next section reviews the standard methodology in more detail and compares it to the additional methodology specifications in the following section.
@@ 3 The Standard Meta-Analysis Methodology
For the above equation $y_{s}$ represents a ($s\times 1$) vector for the dependent variable, $\mathbf{x}_{s}^{\top}=(1,x_{s1},x_{s2},...,,x_{sp})$ represents a ($s\times p$) vector of covariates per study $s$ (which can include control variables but not fixed effect controls), $\boldsymbol{\beta}=(\beta_{0},\beta_{1},...,\beta_{p})$ represents a ($p\times 1$) vector of effect sizes (i.e. regression coefficients), and $\varepsilon_{s}$ represents a ($s\times 1$) vector for the error term. Note that data used in meta-analyses or meta-regressions can be either at the subject level or aggregate level (Cooper & Patall, 2009; Pereira et al., 2010; Gambarota & Altoè, 2024; Viechtbauer, 2007; Poom & Af Wåhlberg, 2022; Yamaguchi et al., 2014). This format will be shared for all the following methodology specifications in this section and the following section. Another way to show study-level fixed or random effects models is through their shared parent model called General Linear Mixed-effects Model (GLMM), this model is shown below in Equation 3.
@@
For GLMM, $y_{s}$ denotes a ($s\times 1$) vector for the dependent variable, $\mathbf{x}_{s}^{\top}=(1,x_{s1},x_{s2},...,,x_{sp})$ representing a ($s\times p$) vector of covariates per study $s$, $\boldsymbol{\beta}=(\beta_{0},\beta_{1},...,\beta_{p})$ representing a ($p\times 1$) vector of regression coefficients (i.e. effect sizes), $\boldsymbol{\alpha}_{s,FE}$ denotes the fixed effects vector, $\boldsymbol{\alpha}_{s,RE}$ denotes the random effects vector, and $\varepsilon_{s}$ denotes the ($s\times 1$) vector for the error term (Viechtbauer, 2007). The fixed effect can be shown as $\boldsymbol{\alpha}_{s,FE}=\gamma_{s}D_{s}$ such that $D_{s}$ is a vector of dummy variables per study and $\gamma_{s}$ is a vector of fixed effects for location $s$. The random effect is drawn from $(0,\sigma^{2}_{\alpha})$ and the error term $\varepsilon_{s}$ is $(0,\sigma^{2}_{\varepsilon})$ distributed. If $\boldsymbol{\alpha}_{s,FE}$ is removed, the model will be equivalent to a random-effects model, while if $\boldsymbol{\alpha}_{s,RE}$ is removed, the model will be equivalent to a fixed-effects model. The mixed effects model is also discussed in this section when both of the fixed and random effects are present. Note that caution should be used if using the mixed effects model, there is simulation literature that found it performed poorly in certain conditions (like location heterogeneity) compared to fixed effects and/or random effects (Bakbergenuly and Kulinskaya, 2018; Kulinskaya et al., 2021; Rodriguez et al., 2023; LeBeau et al., 2018). Note that we show both the typical syntax of the meta-regression simulation literature as well as the vector notation commonly used in the econometrics literature. The remaining equations will only show the syntax of the meta-regression simulation literature.
@@
Appendix G shows the following details for each methodology specification: the meta-regression framework from statistics/econometrics, parameter estimation, and standard error estimation. Note that all methodology specifications are very similar grouped at the study-level, but their differences lie between whether they use fixed, random and mixed effects. For all the methods in Section 3 & 4, the parameter is estimated with the weighted least square (WLS) estimation strategy and the standard error is estimated with a weighted inverse-variance. WLS is used instead of ordinary least squares (OLS) to provide weights for each group (in this case each study), or the omitted variable bias would likely occur (due to the omitted study-level heterogeneity). This omitted variable bias would then harm performance since it makes the estimator biased (Wooldridge, 2010). Note that there are some studies in the literature that suggest using WLS as both the estimation strategy and model specification, but the focus of this manuscript is a WLS estimation strategy with various random or fixed effects model specifications (T. D. Stanley & Doucouliagos, 2017; Van Aert & Jackson, 2019; T. Stanley et al., 2023). Both approaches are designed to address heterogeneity of meta-analysis data sets, the typical form of the WLS estimator is shown below in Equation 4.
@@
4 Additional Meta-Analysis Methodologies:
Modeling Location & Time
In the following subsections, we briefly discuss methodology specifications that control for location and/or time variation: location-level random/fixed/mixed effects are presented in Section 4.1, time-level fixed effects are presented in Section 4.2, combined location and time-level fixed effects are presented in Section 4.3, and specifications that model time with a trend combined with study-level and location-level fixed effects are presented in Section 4.4. Location and study-level fixed effects with a trend term have been introduced in the literature, and the other specifications have been added to ensure that all relevant variations of the methodology specification are tested. These effects help to account for differences across groups (i.e. geographic locations) or across time periods that may otherwise bias the estimates of the treatment effect. By including these effects, we can obtain more accurate and unbiased estimates of the relationships between covariates and effect sizes by accounting for heterogeneity with precision beyond study-level. Note that there are only a few random and mixed effect specifications since our simulation is designed so the random effects assumption does not hold (since it is rare in practice). All of these methodologies will be compared to the standard meta-regression methodology, and then all of these methods will be tested in the simulation in the next section with data that have joint location and time heterogeneity between the studies in a meta-regression. The performance criteria in Section 5.2 will dictate which methodology specification performs best in each situation. Hypothesis 1 will be tested by comparing two joint heterogeneity methodology specifications with two non-joint heterogeneity methodology specifications. Hypothesis 2 will be tested by comparing the standard method to the joint heterogeneity methodology specifications. Hypothesis 3 will be tested with all of the non-standard specifications.

### v2

@@ Abstract
In this paper, we conduct a simulation study with subject level data to evaluate conventional meta-regression approaches (study-level random, fixed, and mixed effects) against seven methodology specifications new to meta-regressions that control joint heterogeneity in location and time (including a new one that we introduce). We systematically vary heterogeneity levels to assess statistical power, estimator bias and model robustness for each methodology specification. This assessment focuses on three aspects: performance under joint heterogeneity in location and time, the effectiveness of our proposed settings incorporating location fixed effects and study-level fixed effects with a time trend, as well as guidelines for model selection. The results show that jointly modeling heterogeneity when heterogeneity is in both dimensions improves performance compared to modeling only one type of heterogeneity.
@@
Competing Disclosure: To the best of our knowledge, no competing interests exist
Funding Statement: This research was not funded
@@ 1 Introduction
A meta-analysis is a useful tool that combines results from independent studies with similar design to examine whether there is consistency in the results across studies or if those results conflict (McGrath et al., 2019; Bakbergenuly and Kulinskaya, 2018; Rice et al., 2018; Kelley, 2012). Meta-regressions are commonly referred to as meta-analyses (and vice versa), since the only difference introduced by a meta-regression is that it transcends the univariate model of a meta-analysis by including identifiable study and/or subject characteristics as covariates. Meta-regression data looks very similar to regression data with covariates, a variable of interest, and group-level controls. The only difference is that the estimators are estimating the common effect sizes after controlling for the study-level heterogeneity, instead of a effect size for a specific data set. Data from the different studies under consideration can be at the subject level or aggregate level using the summary statistics of each study; this manuscript will focus on the subject level due to the many issues inherent in aggregate level data (Yamaguchi et al., 2014; Cooper & Patall, 2009; Lambert et al., 2002; Stewart & Clarke, 1995). Consequently, thousands of meta-analyses have been performed, primarily in the medical, pharmaceutical, ecology, and social sciences literature (Borenstein et al., 2009). The methodology used for meta-analyses is designed to accurately estimate the effect sizes and mitigate any heterogeneity that may be present between each of the studies in the meta-analysis. The standard methodology uses study-level effects: most commonly random effects, fixed effects, and on rare occasions mixed effects (Gambarota & Altoè, 2024; T. Stanley et al., 2023; Sugasawa & Noma, 2021; Richardson et al., 2021; Pateras et al., 2018; T. D. Stanley & Doucouliagos, 2017). But what about heterogeneity in location or time? It is common for meta-analyses to include decades worth of studies from many locations around the world, which introduces the likely variation in both the time and location in which each study was conducted (Thöni & Volk, 2021; Van Den Akker et al., 2020; Engel, 2011; Oosterbeek et al., 2004). In the experimental economic and meta-analysis simulation literature, almost all meta-analyses are conducted at the study level (dominantly using random effects), with only a small percentage controlling for location variation and none controlling for time variation. [^1] Table 4 in Appendix A shows a representative sample of meta-regressions in experimental economics and how they control for heterogeneity: only about 30% control for location heterogeneity, only about 15% control for precise country-level heterogeneity, and only about 5% control for time heterogeneity (implicitly). This manuscript tests three hypotheses when the data follows the standard assumptions:
@@
This research aims to test which meta-regression methodology performs optimally when joint heterogeneity in the location and time between studies in a meta-regression is present in varying degrees. To achieve our research objective, we conduct a simulation study that tests the performance of ten meta-regression methodology specifications that include the standard methodology and the specifications that control for heterogeneity in location and/or time. Note that most of the methodology specifications are fixed-effect specifications since the assumption of random effects is a phenomenon that rarely holds in practice (which is required for random or mixed effect models, please see Section 5.1 for further details) (Clark & Linzer, 2015; Wooldridge, 2010; Baltagi, 2005; Hausman & Taylor, 1981; Mundlak, 1978). In summary, our simulation found that the standard methodology performed the best and the specifications that control for joint heterogeneity performed a close second best. This excludes study-level fixed effects with a trend proposed in the meta-analysis simulation literature because this provided inaccurate trend estimates, but includes a new specification inspired by the literature and introduced by this manuscript: location-level fixed effects with a trend. Accordingly, specifications that did not jointly control for location and time heterogeneity performed the worst (including location-level fixed effects proposed in the experimental economics literature). Note that these results are for joint heterogeneity, in future research we plan to test when only location-level heterogeneity is present. The organization of this paper is as follows: Section 2 reviews the relevant literature, Section 3 shares the methodologies of a standard meta-analysis, and Section 4 shows the additional methodology specifications that control for location and/or time heterogeneity. The simulation in Section 5 tests the performance of each of the methods in the following way: simulation design, the performance criteria of each specification, and which specification performs optimally. An essential step-by-step practitioners guide is outlined in Section 6. Section 7 shows the utility of example applications as well as how simulations can be extended beyond purely synthetic data by extracting parameters from real data. Lastly, Section 8 provides concluding remarks and notes on the related areas we plan to study in the future.
@@ 2 Literature Review
There is dense literature that argues whether meta-analyses and meta-regressions should utilize a study-level fixed effects model or a study-level random effects model (Gambarota & Altoè, 2024; Sugasawa & Noma, 2021; Viechtbauer, 2007; Brockwell & Gordon, 2001; Hunter & Schmidt, 2000). Random effects are the most dominant because their design handles data randomly drawn from a larger population (such as experimental subject data), can capture the sampling error of the participants, and can also capture the study-level variance of the participant sampling for each experiment (Kelley, 2012; Higgins et al., 2009; Cooper & Patall, 2009). However, a random effects model requires that variation in draws of the random effect are uncorrelated with variation in all other explanatory variables, which is rare in practice (Clark & Linzer, 2015; Wooldridge, 2010; Baltagi, 2005; Hausman & Taylor, 1981; Mundlak, 1978). The random effects model should not be used when this assumption is violated, the consequences of this model misspecification can be most easily illustrated in the inflated estimator bias (Van Aert & Jackson, 2019; Spanos, 2019b). There is also a less common mixed effects model (also known as the generalized linear mixed model). This model combines fixed and random effects to capture both the variation within each group (with fixed effects) as well as the variation across groups (with random effects); mixed effects are commonly used for hierarchical data where data can be categorized into groups across different levels (Papadimitropoulou et al., 2019; Bakbergenuly and Kulinskaya, 2018; Stijnen et al., 2010; Viechtbauer, 2010; Overton, 1998). There are important trade-offs to consider for all of these methodology specifications. For instance, one study found an overestimation of standard errors with random effects and an underestimation of standard errors with fixed effects (Overton, 1998). The tradeoffs are compounded for mixed effects since it includes the benefits of both fixed and random effects, but also the limitations of both fixed and random effects. This is related to our research because this is the standard methodology that is tested and compared to our methodology specifications that control location and/or time. To recap, do not indiscriminately use random effects; it has assumptions and limitations of their own, just like fixed effects and mixed effects.
@@
This literature also covers methodology performance when varying degrees of study-level heterogeneity are present (McGrath et al., 2019; Bakbergenuly and Kulinskaya, 2018; Sánchez-Meca & Marín-Martínez, 2008; Hedges & Vevea, 1998; Biggerstaff and Tweedie, 1997). The common finding is that when higher heterogeneity between studies is present (compared to lower heterogeneity), neither study-level random or fixed effects perform as well. Our research extends this to varying degrees of heterogeneity at the location and time level. This is important since the literature overlooks the possibility of variation in location or time in a meta-analysis (excluding only a select few studies). The only exception are spatiotemporal meta-analyses, but the focus of this field of literature is very different since these use spatial economic models to model the heterogeneity between locations by using the distance between them along with other related covariates (B. T. Johnson et al., 2017). In contrast, when this manuscript describes location heterogeneity, its focus is closer to two way fixed effects in panel data models: distance apart is not a factor, we solely control whether locations are different or not. Note that many of the sources that highlight spatiotemporal heterogeneity is present, don’t account for it explicitly in their modeling (Jiang et al., 2024; Liu & Xiao, 2022; Ren et al., 2022). Recently, some of those studies have explored controlling for location-level or time-level effects more directly than implicitly through study-level effects. In the experimental economic literature, location-level effects were controlled by country, geographic region, or continent (Aimone et al., 2024; N. D. Johnson & Mislin, 2011; Oosterbeek et al., 2004). Our research analyzes the performance of each specification and how well their specification captures varying degrees of location-level heterogeneity, neither of these have been done since the aforementioned meta-analyses are only empiric.
@@
The meta-analysis simulation literature explores time-level effects, but each study controls or measures time differently, so there is no set consensus. Accordingly, our research aims to set a consensus by testing the relevant specifications that control for time heterogeneity. Two studies measure fixed effects with a trend for experiments that take place at different time periods, and another study introduces a meta-analysis methodology called Continuous Time Meta-Analysis (CoTiMA) that includes time lags for both within each study and between the studies inside of a meta-analysis (LeBlanc & Banks, 2022; Rudolph et al., 2020; Dormann et al., 2020). Closest to our research is a study that introduces a meta-analysis methodology to model variation over time between the static studies inside the meta-analysis (LeBlanc & Banks, 2022). This methodology models variation in time by testing if there are any trends in each static study across time by category. So, only the date in the meta-analysis is dynamic; there are no time lags within each study. But, either time-level fixed effects or fixed effects with a trend capture the variation of time between studies (please see Section 4 for further details). Both studies used study-level fixed effects with a trend, so we are going to test this specification in the face of joint location and time heterogeneity and compare its performance to other specifications. These other specifications include the standard methodology, location-level fixed effects with a trend, and others that control for location and/or time. Our manuscript is the first to test the following meta-regression specifications together in a simulation: location-level fixed/random effects, time-level fixed effects, and location-level fixed effects with trend. The goal of testing these methods are different from spatiotemporal heterogeneity meta-analyses because they control for the heterogeneity instead of spatially modeling it. The next section reviews the standard methodology in more detail and compares it to the additional methodology specifications in the following section.
@@ 3 The Standard Meta-Analysis Methodology
For the above equation $y_{s}$ represents a ($s\times 1$) vector for the dependent variable, $\mathbf{x}_{s}^{\top}=(1,x_{s1},x_{s2},...,,x_{sp})$ represents a ($s\times p$) vector of covariates per study $s$ (which can include control variables but not fixed effect controls), $\boldsymbol{\beta}=(\beta_{0},\beta_{1},...,\beta_{p})$ represents a ($p\times 1$) vector of effect sizes (i.e. regression coefficients), and $\varepsilon_{s}$ represents a ($s\times 1$) vector for the error term. Note that data used in meta-analyses or meta-regressions can be either at the subject level or aggregate level (Gambarota & Altoè, 2024; Poom & Af Wåhlberg, 2022; Yamaguchi et al., 2014; Pereira et al., 2010; Viechtbauer, 2007; Cooper & Patall, 2009). This format will be shared for all the following methodology specifications in this section and the following section. Another way to show study-level fixed or random effects models is through their shared parent model called General Linear Mixed-effects Model (GLMM), this model is shown below in Equation 3.
@@
For GLMM, $y_{s}$ denotes a ($s\times 1$) vector for the dependent variable, $\mathbf{x}_{s}^{\top}=(1,x_{s1},x_{s2},...,,x_{sp})$ representing a ($s\times p$) vector of covariates per study $s$, $\boldsymbol{\beta}=(\beta_{0},\beta_{1},...,\beta_{p})$ representing a ($p\times 1$) vector of regression coefficients (i.e. effect sizes), $\boldsymbol{\alpha}_{s,FE}$ denotes the fixed effects vector, $\boldsymbol{\alpha}_{s,RE}$ denotes the random effects vector, and $\varepsilon_{s}$ denotes the ($s\times 1$) vector for the error term (Viechtbauer, 2007). The fixed effect can be shown as $\boldsymbol{\alpha}_{s,FE}=\gamma_{s}D_{s}$ such that $D_{s}$ is a vector of dummy variables per study and $\gamma_{s}$ is a vector of fixed effects. The random effect is drawn from $N(0,\sigma^{2}_{\alpha})$ and the error term $\varepsilon_{s}$ is $N(0,\sigma^{2}_{\varepsilon})$ distributed. If $\boldsymbol{\alpha}_{s,FE}$ is removed, the model will be equivalent to a random-effects model, while if $\boldsymbol{\alpha}_{s,RE}$ is removed, the model will be equivalent to a fixed-effects model. The mixed effects model is also discussed in this section when both of the fixed and random effects are present. Note that caution should be used if using the mixed effects model, there is simulation literature that found it performed poorly in certain conditions (like location heterogeneity) compared to fixed effects and/or random effects (Rodriguez et al., 2023; Kulinskaya et al., 2021; LeBeau et al., 2018; Bakbergenuly and Kulinskaya, 2018). Note that we show both the typical syntax of the meta-regression simulation literature as well as the vector notation commonly used in the econometrics literature. The remaining equations will only show the syntax of the meta-regression simulation literature.
@@
Appendix F shows the following details for each methodology specification: the meta-regression framework from statistics/econometrics, parameter estimation, and standard error estimation. Note that all methodology specifications are very similar grouped at the study-level, but their differences lie between whether they use fixed, random and mixed effects. For all the methods in Section 3 & 4, the parameter is estimated with the weighted least square (WLS) estimation strategy and the standard error is estimated with a weighted inverse-variance. WLS is used instead of ordinary least squares (OLS) to provide weights for each group (in this case each study), or the omitted variable bias would likely occur (due to the omitted study-level heterogeneity). This omitted variable bias would then harm performance since it makes the estimator biased (Wooldridge, 2010; Cameron & Trivedi, 2005; Baltagi, 2005). Note that there are some studies in the literature that suggest using WLS as both the estimation strategy and model specification, but the focus of this manuscript is a WLS estimation strategy with various random or fixed effects model specifications (T. Stanley et al., 2023; Van Aert & Jackson, 2019; T. D. Stanley & Doucouliagos, 2017). Both approaches are designed to address heterogeneity of meta-analysis data sets, the typical form of the WLS estimator is shown below in Equation 4.
@@
4 Additional Meta-Analysis Methodologies: Location & Time
In the following subsections, we briefly discuss methodology specifications that control for location and/or time variation: location-level random/fixed/mixed effects are presented in Section 4.1, time-level fixed effects are presented in Section 4.2, combined location and time-level fixed effects are presented in Section 4.3, and specifications that model time with a trend combined with study-level and location-level fixed effects are presented in Section 4.4. Location and study-level fixed effects with a trend term have been introduced in the literature, and the other specifications have been added to ensure that all relevant variations of the methodology specification are tested. These effects help to account for differences across groups (i.e. geographic locations) or across time periods that may otherwise bias the estimates of the treatment effect. By including these effects, we can obtain more accurate and unbiased estimates of the relationships between covariates and effect sizes by accounting for heterogeneity with precision beyond study-level. Note that there are only a few random and mixed effect specifications since our simulation is designed so the random effects assumption does not hold (since it is rare in practice). All of these methodologies will be compared to the standard meta-regression methodology, and then all of these methods will be tested in the simulation in the next section with data that have joint location and time heterogeneity between the studies in a meta-regression. The performance criteria in Section 5.2 will dictate which methodology specification performs best in each situation. Hypothesis 1 will be tested by comparing two joint heterogeneity methodology specifications with two non-joint heterogeneity methodology specifications. Hypothesis 2 will be tested by comparing the standard method to the joint heterogeneity methodology specifications. Hypothesis 3 will be tested with all of the non-standard specifications. Appendix G shows the following details for each methodology specification in this section: the meta-regression framework from statistics/econometrics, parameter estimation, and standard error estimation.